OpenStudy (anonymous):

2=3x+22/5 Solve for x

OpenStudy (anonymous):

Is it x= -4/5 ?

OpenStudy (anonymous):

. Yes! you are right!

OpenStudy (campbell_st):

is the equation \[2 = \frac{3x +22}{5}\] or \[2 = 3x + \frac{22}{5}\]

OpenStudy (anonymous):

it was the firdt one

OpenStudy (anonymous):

Urch, your answer is right

OpenStudy (campbell_st):

multiply both sides by 5 then 10 = 3x + 22 subtract 22 -12 = 3x divide by 3 x = -4